Find the sum of binomial coefficients: (30)+(40)+(40)+(20...+20
Answer #1

C c² + ......... t och (1+x) = cot tin x t al but X=1 and nz 10 jogo (1 +0930 = 10 + 10 + 100+ . 3 10)+(10)+1)+ GS+.. . + ()